# Heights and periods of algebraic cycles in families

### v1

@@ Abstract.
We consider the Beilinson–Bloch heights and Abel–Jacobian periods of homologically trivial Chow cycles in families. For the Beilinson–Bloch heights, we show that for any $g\geq 3$, there is a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over $\overline{\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over $U$ satisfy the Northcott property. For the Abel–Jacobi periods, we provide an algebraic criterion for the existence of a Zariski open dense subset of any family such that all cycles not defined over $\overline{\mathbb{Q}}$ are non-torsion and verify that this criterion holds true for Ceresa cycles and Gross–Schoen cycles.
@@ 1. Introduction
This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in families. We have achieved two goals: the Northcott property for the heights of the Gross–Schoen cycles and Cerasa cycles and an algebraic criterion for the non-degeneracy of Abel–Jacobi periods for the general family of cycles.
1.1. Gross–Schoen and Ceresa cycles
For a smooth projective irreducible curve $C$ defined over a filed $k$ of $g\geq 3$ and a class $\xi\in\mathrm{Pic}^{1}(C)$ such that $(2g-2)\xi=\omega_{C}$, we have two homologically trivial 1-cycles:
@@ 1.1. Gross–Schoen and Ceresa cycles
    the Gross–Schoen cycle ${\mathrm{GS}}(C):=\Delta_{\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
@@
    the Ceresa cycle by ${\mathrm{Ce}}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$.
Up to torsions, the definition of these cycles does not depend on the choice of $\xi$.
Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
@@
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$, and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$,
$\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
$\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$.
@@
For each $g\geq 3$, there exists a Zariski open dense subset of the ${\mathscr{M}}_{g}$ defined over $\overline{\mathbb{Q}}$ such that for any $H,D\in{\mathbb{R}}$,
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty,$
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty.$
In the course of the proof, we also establish the following geometric result.
@@
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ such that the followings hold true:
@@
    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in U({\mathbb{C}})\setminus U(\overline{\mathbb{Q}})$.
@@
    there exist at most countably many $s\in U({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
Hain [Ha24] also proved Theorem 1.3 with $U$ a non-empty analytic open subset of ${\mathscr{M}}_{g}^{\mathrm{an}}$ with a different method. When $g=3$, Hain’s result is completely explicit, while ours is not.
The proofs of Theorem 1.1, Corollary 1.2 and Theorem 1.3 have two parts: the arithmetic part and the geometric part. Both parts are needed to prove Theorem 1.1 and Corollary 1.2, while Theorem 1.3 only relies on the geometric part. We will explain each part in more detail in §$\S$1.3.
Before moving on, let us mention that the geometric part of our paper works for any family of homologically trivial cycles and, more generally, for any admissible normal function. More specifically, for any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S_{{\mathscr{F}}}(1)$ of $S$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S_{{\mathscr{F}}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian. We then prove a checkable criterion for $S_{{\mathscr{F}}}(1)\not=S$. A simple case will be presented in Corollary 1.9 and Remark 1.10. We do these by studying the normal function associated with $Z$, which is defined in §$\S$§C.5, and by studying its Betti rank.
1.2. Normal functions
Now, we turn to the geometric part of the framework and explain our results.
Let $S$ be a quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$, and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
@@ 1.2. Normal functions
Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$.
@@
Let ${\mathscr{F}}_{\mathrm{Betti}}$ denote the induced foliation, which we call the Betti foliation. More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $\Delta$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a flat section of ${\mathbb{V}}_{\mathbb{R}}$. The manifolds $\sigma(U)$ gluing together to a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. See §$\S$§C.2 for more details. In particular, all torsion multi-sections are leaves of ${\mathscr{F}}_{\mathrm{Betti}}$.

### v2

@@ Abstract.
We consider the Beilinson–Bloch heights and Abel–Jacobian periods of homologically trivial Chow cycles in families. For the Beilinson–Bloch heights, we show that for any $g\geq 3$, we can construct a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over ${\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over ${\mathscr{M}}_{g}^{\mathrm{amp}}$ have a lower bound and satisfy the Northcott property. For the Abel–Jacobi periods, we prove that the Ceresa and Gross–Schoen cycle over every $\overline{\mathbb{Q}}$ point in ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is non-torsion.
We also prove results and make conjectures for any family $X\rightarrow S$ of smooth projective varieties and homologically trivial cocycles $Z$ defined over a number field. Using the Betti strata, we construct a Zariski open subset $S^{\mathrm{amp}}$ of $S_{{\mathbb{C}}}$ over which $Z_{s}$ is non-torsion for all transcendental points $s$ and prove a criterion for $S^{\mathrm{amp}}$ to be dense. We conjecture the rationality of $S^{\mathrm{amp}}$ and also conjecture that $S^{\mathrm{amp}}$ is precisely the open locus on which the Beilinson–Bloch height of $Z_{s}$ has a lower bound and satisfies the Northcott property.
Dedicated to Gerd Faltings’s 70th birthday
@@ 1. Introduction
For a curve over ${\mathbb{C}}$, the classical Abel–Jacobi map is a bijection
${\mathrm{AJ}}\colon\mathrm{Pic}^{0}(C){\overset{\sim}{{\longrightarrow}}}{\mathrm{Jac}}(C).$
When $C$ is defined over a number field $k$, one can define the Néron–Tate height on ${\mathrm{Jac}}(C)(\bar{k})$ using the canonical polarization. This height has a Northcott property, which gives positivity of the height pairing, which in turn is an important ingredient for the proof of the Mordell–Weil theorem and the formulation of the Birch and Swinnerton-Dyer conjecture. Notice that by the Hodge index theorem of Faltings and Hriljac, these heights can also be defined as arithmetic intersections on integral models of $C$.
In the 1950s-1960s, Weil and Griffiths extended the Abel–Jacobi map to homologically trivial cycles (say of codimension $n$) on smooth and projective varieties $X$ over ${\mathbb{C}}$
${\mathrm{AJ}}^{n}\colon{\mathrm{Ch}}^{n}(X)_{\mathrm{hom}}{\longrightarrow}{\mathrm{Jac}}^{n}(X)$
where ${\mathrm{Jac}}^{n}(X)$ are called intermediate Jacobians and are only complex tori in general. In the 1980s, Beilinson [Blo84] and Bloch [Blo84] independently proposed a conditional definition of heights for these cycles using integral models. They conjectured the positivity of their heights, an extension of the Mordell–Weil theorem, and an extension of the Birch and Swinnerton-Dyer conjecture. However, the positivity cannot be proved via a Northcott property because of the lack of algebraicity of ${\mathrm{Jac}}^{n}(X)$.
This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in a family of smooth and projective varieties $f\colon X\rightarrow S$ parametrized by a quasi-projective variety $S$. We have achieved two goals:
@@
    The generic positivity and Northcott property for the Beilinson–Bloch heights of the Gross–Schoen cycles and Cerasa cycles parametrized by moduli of curves. These are the simplest situations where the Beilinson–Bloch heights are unconditionally defined.
@@
    An algebraic criterion for the non-degeneracy of Abel–Jacobi periods for general families of cycles.
Our method of proof, roughly speaking, is to utilize the intermediate Jacobians to get some geometric positivity and then extend it to some arithmetic positivity. In the following, we describe the main results and the methods of proof in detail.
1.1. Main results on Gross–Schoen and Ceresa cycles
For a smooth projective irreducible curve $C$ defined over a field $k$ and a class $\xi\in\mathrm{Pic}^{1}(C)$, we have two homologically trivial 1-cycles:
  • the Gross–Schoen cycle based at $\xi$ defined by ${\mathrm{GS}}_{\xi}(C):=\Delta_{{\mathrm{GS}},\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
  • the Ceresa cycle based at $\xi$ defined by ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$ ([Cer83]).
Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$, let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
For each $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})({\mathbb{C}})$, to ease notation denote by
${\mathrm{GS}}(\xi):={\mathrm{GS}}_{\xi}({\mathscr{C}}_{s})\quad\text{ and }\quad{\mathrm{Ce}}(\xi):={\mathrm{Ce}}_{\xi}({\mathscr{C}}_{s})$
for $s=\pi(\xi)\in{\mathscr{M}}_{g}({\mathbb{C}})$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
Here are our main results about the Beilinson–Bloch heights for these cycles unconditionally defined by Gross–Schoen [GS95] and Künnemann [Kün01] respectively.
@@ 1.1. Main results on Gross–Schoen and Ceresa cycles
For each $g\geq 3$, there exist a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$ and $h_{\mathrm{NT}}$ is the canonical Néron–Tate height on ${\mathrm{Jac}}({\mathscr{C}}_{s})(\overline{\mathbb{Q}})$.
The proof will be executed in §$\S$6.3, in which we give a construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$. Indeed, we will take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, where ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is the Betti degeneracy locus (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6). A priori ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is only a real analytic set. But we show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is proper Zariski closed in ${\mathscr{M}}_{g,{\mathbb{C}}}$ by applying Ax-Schanuel multiple times; see Theorem 1.4 and Corollary 1.8. Finally, we show that it is defined over ${\mathbb{Q}}$ by volume calculation for adelic line bundles; see Corollary 6.4. Because of the importance of this Zariski open dense set ${\mathscr{M}}_{g}^{\mathrm{amp}}$ on which (1.1) holds, we will make the following definition: We call
  1. (i)
    ${\mathscr{M}}_{g}^{\mathrm{amp}}$ the ample locus of the Gross–Schoen and Ceresa cycles;
  2. (ii)
    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles. By the explanation above, it is the ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. It certainly contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm. C].
We expect ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is the largest open subset of ${\mathscr{M}}_{g}$ so that the height inequality (1.1) in Theorem 1.1 holds; see Conjecture 1.12.
@@
For each $g\geq 3$ and the Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of the ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ as above, we have the following Northcott property: For any $H,D\in{\mathbb{R}}$,
#$\#\{\xi\in\pi^{-1}({\mathscr{M}}_{g}^{\mathrm{amp}})(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(\xi):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}<H\}<\infty,$
#$\#\{\xi\in\pi^{-1}({\mathscr{M}}_{g}^{\mathrm{amp}})(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(\xi):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}<H\}<\infty.$
In the course of the proof, we also establish the following geometric result on the torsion locus. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{s})$.
@@
For each $g\geq 3$ and the Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of the ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ as above, the followings hold true:
@@
    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}({\mathbb{C}})\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$.
@@
    there exist at most countably many $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
Independently, Hain [Hai24] and Kerr–Tayou [KT24] proved Theorem 1.3 with different methods, with ${\mathscr{M}}_{g}^{\mathrm{amp}}$ replaced by a non-empty open subset $U$ of ${\mathscr{M}}_{g,{\mathbb{C}}}$ which is real analytic and complex algebraic respectively.
The proofs of Theorem 1.1, Corollary 1.2, and Theorem 1.3 have two parts: the arithmetic part and the geometric part. Both parts are needed to prove Theorem 1.1, Corollary 1.2, and Theorem 1.3. We will explain each part in more detail in §$\S$1.4.
Before moving on, let us mention that the geometric part of our paper works for any family of homologically trivial cycles and, more generally, for any admissible normal function. We prove a formula for its Betti rank and prove the Zariski closedness of the Betti strata. A simple application is the following result on the torsion locus: For any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S^{\mathrm{Betti}}(1)$ of $S_{{\mathbb{C}}}$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S^{\mathrm{Betti}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S^{\mathrm{Betti}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian. We then prove a checkable criterion for $S^{\mathrm{Betti}}(1)\not=S$. A simple case will be presented in Corollary 1.9 and Remark 1.10. In the case of Gross–Schoen or Ceresa cycles, we will show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$.
1.2. Normal functions, Betti strata and Betti rank
Now, we turn to the geometric part of the framework and explain our results. Let $S$ be a quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$,[1][1][1]Any VHS of odd weight can be reduced to weight $-1$ by applying a suitable twist; see §$\S$§C.1. and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
@@ 1.2. Normal functions, Betti strata and Betti rank
Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures (VMHS) ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$. To any family of homologically trivial cycles, we can associate such an admissible normal function; see §$\S$§C.5.
@@
Let ${\mathscr{F}}_{\mathrm{Betti}}$ denote the induced foliation, which we call the Betti foliation. More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $U$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a section of the local system ${\mathbb{V}}_{\mathbb{R}}$. The manifolds $\sigma(U)$ gluing together to a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. See §$\S$§C.2 for more details. In particular, all torsion multi-sections are leaves of ${\mathscr{F}}_{\mathrm{Betti}}$.